OpenStudy (beth12345):

how to solve |x+4|=|x-2|

OpenStudy (anonymous):

you have two equations to solve \[x+4=x-2\] which clearly has no solution, and \[x+4=2-x\] which does have a solution

OpenStudy (beth12345):

ok